# Complexity at Scale: A Quantitative Analysis of an Alibaba Microservice Deployment

### v2

@@ 4.7. Call Graph Variability
4.8. Heterogeneity Implications
MSs can be multiplexed by a large variety of other MSs and front-end functionality: a common idea employed by many fault management approaches is that the behaviour of MSs is naturally correlated with the usage of front-end applications and their KPIs, and thus deviations in this relationship can be used to detect anomalies (Samir and Pahl, 2019, 2019) and correlation with faulty KPIs can help localise the root cause of faults (Wang et al., 2018; Ma et al., 2019, 2021; Li et al., 2021; Liu et al., 2021). However, we found that MSs can be involved with potentially thousands of front-end service functionalities, inevitably resulting in signal mixing as the behaviour of MSs is the aggregate of many different front-end usage patterns. Without appropriate treatment of signal mixing, for instance through blind source separation (Pal et al., 2013), these approaches are likely to incorrectly detect and attribute faults.
Another approach to anomaly detection is to use statistical thresholds, such as the use of the 3-Sigma rule (Lin et al., 2018a; Yu et al., 2021; Li et al., 2022; Lee et al., 2023) or the extreme studentized deviate test (Vallis et al., 2014), to determine anomalous metric data points. However, such approaches often work under the assumption of approximately normally distributed data. Instead, our findings suggest that MSs can be multiplexed by many front-end workloads, potentially producing non-Gaussian mixtures with varying means and variances, which can yield unexpected tail behaviours and non-stationary signals as different workloads dominate over time. These characteristics could lead to increased false alarms and decreased detection accuracy from such approaches.
One set of popular approaches for leveraging distributed tracing data for anomaly detection and localisation treat the calls within traces as sequences of events (Nedelkoski et al., 2019; Meng et al., 2020; Purfallah Mazraemolla and Rasoolzadegan, 2024). Sequence-learning approaches, such as recurrent neural networks, can then be used to learn the normal sequences of calls between MSs. Trained models can then be used to flag anomalous calls that fall below some threshold of expected next events. However, we found that MSs can have very large out-degrees, meaning a vast number of possible next events. This can dramatically increase the amount of training data required for such approaches. Additionally, a dispersed probability distribution over many potential next events could make approaches sensitive to threshold parameters. Furthermore, picking a one-size-fits-all threshold when the number of next events across MSs is heavy-tailed is likely to result in increased false negatives and increased false positives for MSs with low and high out-degree respectively.
Large‑scale MSAs generate many call graphs, with each front‑end function following a common structure but sometimes producing rare, divergent ones: another common approach employed for fault detection and localisation approaches is to learn the structure of normal system traces. One set of approaches leverage structural modelling to construct propagation graphs based on the structure of normal tracing data (Yang et al., 2023; Panahandeh et al., 2024). However, rare call graphs may be missing from historical data used to construct these models, especially when considering the trace sampling commonly employed in large-scale MSAs, leading to incomplete models and unsatisfactory false positives at runtime. An alternative approach is to use distance-based clustering to group historical traces based on their structure (Wang et al., 2020; Meng et al., 2021). At runtime, new traces are compared against these clusters and any trace that does not match above a given threshold are flagged as anomalous. However, we identified 79 million unique call graphs, making a pair-wise structural distance calculation computationally prohibitive. We found that unique call graphs, even when from the same front-end functionality, can be dissimilar, meaning that this approach could produce a large number of clusters. As each new trace must be compared to the centroid of these clusters, and we found 44 million front-end invocations per hour (Section 3.3), the computational overhead of this approach may be prohibitive at runtime. Furthermore, similar to propagation graph approaches, rare traces may not be captured in historical data, and thus lead to unsatisfactory false positives.
The spread of workload across MSs is highly uneven: our findings, Section 4.3, which corroborate previous findings (Luo et al., 2021, 2022), suggest that focusing on more accurate workload prediction for a small subset of high-demand MSs could yield higher than expected returns in resource savings and SLA compliance. Conversely, the vast majority of MSs, which receive very few calls, may not justify the cost of developing and maintaining sophisticated predictive scaling algorithms. This may be particularly true for MSs that receive such a low amount of workload that they are unlikely to need to scale at all, or could do so efficiently based on more simplistic heuristic rules. Overall, this invites future work on more tailored auto-scaling strategies that focus on optimising prediction accuracy where it matters most.
MS instances are deployed across a large number of nodes and can be co-located with a large number of different MS instances: many approaches for anomaly detection and fault localisation within MSAs leverage distributed tracing data to capture the direct relationships between MSs (Liu et al., 2021; Yang et al., 2023; Panahandeh et al., 2024; Yu et al., 2021, 2023; Gan et al., 2021, 2024). However, in real-world MSAs, MS instances are co-located on the same host to minimise resource underutilisation. Therefore, faults can propagate indirectly between co-located MSs, that do not necessarily share a direct calling relationship, due to resource interference and contention (Ma et al., 2020). Our findings indicate that MSs can be exposed to large numbers of co-located MSs (See Section 4.2). This exposes them to a wide array of possible sources of such fault propagation. As approaches that solely leverage distributed traces cannot capture these faults, future research must be carried out to explore their prevalence and develop approaches that address indirect fault propagation.
5. Dynamicity
MSAs are inherently dynamic. Whilst dependencies are fixed in underlying code, at runtime, calls between MSs are transient, only occurring based on a front-end request. Furthermore, the deployment of MS instances is elastic, with instances being scaled up and down in response to changes in demand and performance. Furthermore, the architecture is prone to change as MSs are added, removed, and updated to evolve the functionality of the system (Newman, 2021). In this section, we explore how this dynamicity manifests in the dataset.
@@ 5.1. Time-Varying Dependencies
A common perspective when discussing MSs within a MSA is the overall dependency structure between MSs (Huye et al., 2023; Gluck, 2020; Gamage and Perera, 2021), as we have done in Section 3.4 4.4. However, this view fails to account for differences in the behaviour of the system at runtime. Indeed, more recent research has shown the value of investigating the temporal behaviour of MSAs at runtime (Parker et al., 2023; Winchester et al., 2023). Still, little research has been carried out to understand to what extent the temporal behaviour of a MSA differs from the aggregate view of the system. As a first step to contributing to this question, we explored the differences in time-varying degree of MSs at runtime by ingesting the call data into the temporal graph database Raphtory (Steer et al., 2024).
@@
Figure 11. Number of deployed instances (blue) and instances normalized by each MS’s maximum number of instances (pink).
Figure 12. Cumulative frequency of the number of times MSs scaled over the 14 days, with lower and upper quartiles (grey dashed) and median (red).
@@ 5.2. Horizontal Scaling
To investigate horizontal scaling within the deployment we used the resource data which includes samples from all deployed MS instances every minute to create a time series, see Figure 11. We found that 76.9% of the 28,000 stateless MS maintained the same number of instances across the 14 days, likely as this dataset does not include all MSs, but only those associated with core services that do not undergo auto-scaling.
MS that scale do not generally do so in response to daily fluctuations in demand: of the remaining 33.1% of MSs we found that only just over half of them (51.2%) scaled more than twice over the 14 days (see Figure 12). To determine the number of MSs that scale daily we extracted both time and frequency domain features using the auto-correlation function (ACF) and fast Fourier transform (FFT). Using the ACF, we derived the ACF ratio by comparing the autocorrelation at a 24-hour lag to that at a lag of 1, providing a measure of the daily cycle’s prominence relative to short-term correlations. Using the FFT derived power spectrum, we calculated the power ratio as a fraction of total spectral power concentrated at the 24-hour frequency. These two features were used as predictors in a Random Forest classifier trained on a manually labelled subset (200 samples) of MSs to identify those exhibiting significant 24-hour periodicity. Using this approach, we identified 1,240 MS that demonstrated strong daily scaling, accounting for only 18.8% of all MSs that scale (i.e., 4.4% of all stateless MS).
MS that scale daily can have a large change in the number of instances: For those MSs that demonstrated strong daily periodicity, we found that the average daily peak to trough change in the number of instances per MS was 13, meaning, on average, MSs experienced a 43.1% change in the number of deployed instances per day.
@@ 5.3. Evolutionary Architecture
The decoupled nature of MSs provides bounded context, meaning that incremental changes can be made to individual MSs without breaking an application, leading to an architecture that grows and changes continuously (Newman, 2021; Huye et al., 2023). As information as to the creation and deprecation of MSs does not exist for this dataset, we assessed the extent to which the deployment undergoes such evolution using the resource data due to its deterministic sampling of deployed MS instances.
@@
Churn occurs on a daily basis: to differentiate between potential scale-to-zero or missing data and true deprecation and creation, we carried out two data processing steps. We first identified every MS with zero instances across the 14 days. Next, if a MS never had instances before or again they were flagged as creation or deprecation candidates respectively. If a candidate had more than one period of zero instances, we calculated the distribution of all period durations, and the pre-creation or post-deprecation period was compared to this distribution. Candidate periods which were greater than 3 standard deviations above the mean were considered as genuine deprecation or creation events.
@@
Results showed that across the 14 days, more MSs were deprecated than created, with 64 new MSs created and 157 MSs deprecated, see Figure 13. Whilst the number of MS creations and deprecations are comparatively small compared to the size of the deployment, the number of new MSs added across two weeks alone was greater than the total size of most experimental testbeds (SocialNetwork - 26 MSs (Delimitrou, 2019), MediaMicroservices - 30 MSs (Delimitrou, 2019), TrainTicket - 68 MSs (FudanSELab, 2018), SockShop - 8 MSs (OCP-Power-Demos, 2023), HipsterShop - 10 MSs (Platform, 2018)).
@@
Alibaba and Meta have similar rates of MS churn: When contrasting the number of MS creation and deprecation events within the dataset to those found by Meta, we found remarkably similar, albeit inverted, values with a median number of MS creations at 5 and 12, and the median number of deprecations at 10.5 and 4 per day for Meta and Alibaba, respectively. The higher deprecation values relative to creation for this dataset when compared to Meta’s may suggest that this deployment, being significantly larger than that of Meta, is in a lifecycle of consolidation and removal of unnecessary MS bloat. However, it should be noted that Meta found significant jumps in creation and deprecation across their longer period of study (Huye et al., 2023) and thus our findings across just 14 days may not accurately depict the broader trend, indeed up until day 6 creation and deprecation events are more in equilibrium. Notably, the median values found for both deployments are of similar magnitudes, despite a considerable difference in the total number of MSs (18,500 vs 28,000), suggesting that the rate of growth of large-scale MSAs may be independent of total size.
@@ 5.4. Dynamicity Implications
The static view of system dependencies does not necessarily correspond to runtime interactions: previous research has used the static view of dependencies to characterise MSAs (Huye et al., 2023) including identifying design anti-patterns (Gamage and Perera, 2021). However, our results in Section 5.1 illustrate that the temporal interactions between MSs can deviate from this static view, with divergent temporal centrality. Therefore, characteristics and attributes identified by applying graph measures to a static view of the MSA (Gamage and Perera, 2021) may produce results that are misaligned with the true behaviour of the system at runtime. Such approaches would likely benefit from pairing the analysis of static representations of the system with temporal, runtime, representation insights afforded by temporal network science (Masuda and Lambiotte, 2016).
MS architectures evolve on a daily basis: of the commonly used MS testbeds (Gan et al., 2019; FudanSELab, 2018; Platform, 2018; OCP-Power-Demos, 2023) only one (Cisco DevNet, 2021) includes versioning that evolves the functionality of the deployment. This may explain why the evolution of MSAs over time is generally not a large consideration of management approaches. As discussed previously, a general approach for anomaly detection and localisation learn the normal behaviour of the system through metric (Yu et al., 2021; Gulenko et al., 2018; Wu et al., 2021, 2020b; Jin et al., 2020; Xu et al., 2017; Wu et al., 2020a; Scheinert et al., 2021) and distributed tracing data (Wang et al., 2020; Yang et al., 2023; Meng et al., 2021; Panahandeh et al., 2024; Nedelkoski et al., 2019; Meng et al., 2020; Purfallah Mazraemolla and Rasoolzadegan, 2024). However, our findings in Section 5.3, in corroboration with Meta’s finding (Huye et al., 2023) and MSA design principles (Newman, 2021), underscore that MSAs are not stationary, with MSs and thus dependencies changing in time. Inevitably, this would lead to the creation and alteration of call graphs, and as MSs are multiplexed by new functionality, the normal behaviour of their metrics can be expected to change. Despite some promising mechanisms to address MSA evolution, such as human-in-the-loop flagging of MSA updates (Purfallah Mazraemolla and Rasoolzadegan, 2024) and transfer learning (Gan et al., 2024), the implication of system evolution on fault management approaches and the additional overhead of frequent retraining is considerably underappreciated in current research.
6. Summary of Research Implications and Testbed Opportunities
@@
In Sections 3.5, 4.8, 5.4 we discussed how management approaches often fail to consider (i) scalability in large-scale deployments (Section 3.5), (ii) the fact that MSs may be multiplexed by many other MSs and front-end functionalities (Section 4.8), (iii) the variability of call graphs from the same front-end functionality (Section 4.8), (iv) the exposure of MSs to potential interference from co-habitation (Section 4.8), and (v) adaptation to an evolving system (Section 5.4).
@@
Many of the above limitations of current state-of-the-art approaches appear to stem, at least in part, from limited access to real-world MS deployments (Huye et al., 2023). The research testbeds that are commonly used throughout the literature (Gan et al., 2019; FudanSELab, 2018; Platform, 2018; OCP-Power-Demos, 2023) are not sophisticated enough to capture the complexities of actual deployments, especially large-scale deployments such as the one studied here or at Meta (Huye et al., 2023). Therefore, we join recent calls for improved testbed realism (Seshagiri et al., 2022; Huye et al., 2023). Our findings allow us to identify several development opportunities that could help testbeds to more accurately represent real-world deployments.
Expanding MSA functionality: MS testbeds are relatively small, containing in the order of 10’s of MSs (Gan et al., 2019; FudanSELab, 2018) or less (OCP-Power-Demos, 2023; Platform, 2018). Furthermore, these deployments only comprise a single application, such as a social network (Gan et al., 2019) or train ticket ordering (FudanSELab, 2018), with limited front-end functionalities. On the other hand, the deployment at Alibaba (Sections 3.2 and 3.1) and Meta (Huye et al., 2023) support many applications and front-end service functionalities. As most testbeds are real small-scale MS deployments, there is significant scope for development to expand their functionality, and subsequently their scale.
Evolving user inputs and architecture: most MS testbeds only provide a single version (Gan et al., 2019; FudanSELab, 2018; Platform, 2018; OCP-Power-Demos, 2023), meaning that there are limited opportunities for the kind of architectural evolution that was found in this dataset (Section 5.3) and at Meta. This underscores the need for continued development of current testbeds, which could naturally facilitate evolutionary behaviours through versioning support. To simulate user behaviour, testbeds use workload generators, however, some of these generators only produce stationary load (Tene, 2012). On the other hand, user demand is in constant flux, as demonstrated in our findings (Section 3.3). The differing usage patterns of distinct front-end services can cause variation in the behaviour of a MSA, such as the prevalence of specific call graphs, and thus dependencies at specific times (Section 5.3, (Winchester et al., 2023)). The processed Alibaba dataset released with this paper (Winchester, 2025) includes the front-end service usage patterns over the 14-days, which can be combined with workload generators (Locust.io, [n. d.]) to create more realistic usage patterns for testbeds.
Call graph variability: testbeds are generally too simplistic in their provided functionality and number of MSs to generate the wide degree of unique call graphs (Section 4.7) and call graph variability (Section 4.7) identified in our investigation and within other MSAs (Luo et al., 2021; Huye et al., 2023). To represent this behaviour, testbeds can be developed to include mechanisms that are known to create variable call graphs such as graceful degradation, which provides alternative call paths under stress or failure conditions, and caching limits that lead to data eviction. Furthermore, front-end functionality can be designed with top-down variability in mind, such as conditional user inputs that trigger alternative call paths.

### v3

@@ 6.1. Time-Varying Dependencies
A common perspective when discussing MSs within a MSA is the overall dependency structure between MSs (huye_lifting_2023; gluck_introducing_2020; gamage_using_2021), as we have done in Section 4.4 and 5.3. However, this view fails to account for differences in the behaviour of the system at runtime. Indeed, more recent research has shown the value of investigating the temporal behaviour of MSAs at runtime (parker_visualizing_2023; winchester_temporal_2023). Still, little research has been carried out to understand to what extent the temporal behaviour of a MSA differs from the aggregate view of the system. As a first step to contributing to this question, we explored the differences in time-varying degree of MSs at runtime by ingesting the call data into the temporal graph database Raphtory (steer_raphtory_2024).
@@
(a) Number of deployed instances (blue) and instances normalized by each MS’s maximum number of instances (pink).
(b) Cumulative frequency of the number of times MSs scaled over the 14 days, with lower and upper quartiles (grey dashed) and median (red).
Figure 11. Dynamicity of MS deployments: (a) scaling patterns across instances, (b) distribution of scaling events.
@@ 6.2. Horizontal Scaling
The traces in this dataset only includes long running MSs which are a part of core applications that do not undergo much auto-scaling (see 3). Nonetheless, to investigate the horizontal scaling within the deployment we used the resource data which includes samples from all deployed MS instances every minute to create a time series, see Figure 11(a). We found that 76.9% of the 28,000 stateless MS maintained the same number of instances across the 14 days.
MS that scale do not generally do so in response to daily fluctuations in demand: of the remaining 33.1% of MSs we found that only just over half of them (51.2%) scaled more than twice over the 14 days (see Figure 11(b)). To determine the number of MSs that scale daily we extracted both time and frequency domain features using the auto-correlation function (ACF) and fast Fourier transform (FFT). Using the ACF, we derived the ACF ratio by comparing the autocorrelation at a 24-hour lag to that at a lag of 1, providing a measure of the daily cycle’s prominence relative to short-term correlations. Using the FFT derived power spectrum, we calculated the power ratio as a fraction of total spectral power concentrated at the 24-hour frequency. Using these two metrics of periodicity we identified 1,240 MS that demonstrated strong daily scaling, accounting for only 18.8% of all MSs that scale (i.e., 4.4% of all stateless MS).
MS that scale daily can have a large change in the number of instances: For those MSs that demonstrated strong daily periodicity, we found that the average daily peak to trough change in the number of instances per MS was 13, meaning, on average, MSs experienced a 43.1% change in their number of deployed instances per day.
@@ 6.3. Evolutionary Architecture
The decoupled nature of MSs provides bounded context, meaning that incremental changes can be made to individual MSs without breaking an application, leading to an architecture that grows and changes continuously (newman_building_2021; huye_lifting_2023). As information as to the creation and deprecation of MSs does not exist for this dataset, we assessed the extent to which the deployment undergoes such evolution using the resource data due to its deterministic sampling of all deployed MS instances every minute.
@@
Churn occurs on a daily basis: to differentiate between potential scale-to-zero or missing data and true deprecation and creation, we carried out two data processing steps. We first identified every MS with zero instances across the 14 days. Next, if a MS never had instances before or again they were flagged as creation or deprecation candidates respectively. If a candidate had more than one period of zero instances, we calculated the distribution of all period durations, and the pre-creation or post-deprecation period was compared to this distribution. Candidate periods which were greater than 3 standard deviations above the mean were considered genuine deprecation or creation events.
@@
Results showed that across the 14 days, more MSs were deprecated than created, with 64 new MSs created and 157 MSs deprecated, see Figure 12. Whilst the number of MS creations and deprecations are comparatively small compared to the size of the deployment, the number of new MSs added across two weeks alone was greater than the total size of most experimental testbeds (SocialNetwork - 26 MSs (noauthor_deathstarbenchsocialnetwork_nodate), MediaMicroservices - 30 MSs (noauthor_deathstarbenchsocialnetwork_nodate), TrainTicket - 68 MSs (noauthor_fudanselabtrain-ticket_2025), SockShop - 8 MSs (noauthor_ocp-power-demossock-shop-demo_2025), HipsterShop - 10 MSs (noauthor_googlecloudplatformmicroservices-demo_2025)).
@@
Alibaba and Meta have similar rates of MS churn: When contrasting the number of MS creation and deprecation events within the dataset to those found by Meta, we found remarkably similar, albeit inverted, values with a median number of MS creations at 5 and 12, and the median number of deprecations at 10.5 and 4 per day for Meta and Alibaba, respectively. The higher deprecation values relative to creation for this dataset when compared to Meta’s may suggest that this deployment, being significantly larger than that of Meta, is in a lifecycle of consolidation and removal of unnecessary MS bloat. However, it should be noted that Meta found significant jumps in creation and deprecation across their longer period of study (huye_lifting_2023) and thus our findings across just 14 days may not accurately depict the broader trend, indeed up until day 6 creation and deprecation events are more in equilibrium. Notably, the median values found for both deployments are of similar magnitudes, despite a considerable difference in the total number of MSs (18,500 vs 28,000), suggesting that the rate of growth of large-scale MSAs may be independent of total size.
Table 4. Common assumptions in MS management approaches, their validity in large-scale MSAs, and recommendations for future work.
Assumption Finding Recommendation
Moderate deployment size. Real deployments can comprise tens of thousands of MSs and millions of instances (4.1). Design and evaluate methods for scalability using large-scale datasets/testbeds (luo_power_2022; lee_tale_2024; detti_ubench_2023).
Availability of labelled data. The enormous diversity of MSs and functionalities (4.1, 4.2) makes creating an adequately representative labelled corpus impractical. Shift toward partial (zhou_wsl_2017), self- (chen_simclr_2020), or synthetic (chawla_smote_2002) supervision.
Per-service modelling is tractable. Thousands of front-end services (4.2) with distinct usage and call graphs makes per-service modelling challenging. Explore shared learning frameworks such as hierarchical or multi-task learning (ruder_multitask_2017) to combine learning across services.
Stationary workloads and call graphs. System-level workload fluctuates across each day (4.3). Evaluate methods under realistic non-stationary workloads (noauthor_locustio_nodate).
Simple coupling between front-end KPIs to single MSs MSs are multiplexed by thousands of front-end functionalities (5.5), likely causing signal mixing. Use latent-factor models (jolliffe_pca_2011) or causal discovery with unobserved confounders (spirtes_cps_2001).
Metric distributions are approximately normal. Multiplexed workloads (5.5) likely produce non-Gaussian mixtures with changing means and variances. Replace Gaussian thresholds with robust heavy-tailed (huber_robust_2009), quantile-based (koenker_quantile_2005), or copula models (nelsen_introduction_2006).
Number of potential next calls in a call graph is moderate. Branching factor in traces can be very large (5.3). Adopt sub-graph pattern mining (yan_gspan_2002) or attention-based sequence models (vaswani_attention_2017).
Call graphs from the same front-end service are stable. Rare, divergent call graphs are common (5.6) Apply graph-level embedding (narayanan_graph2vec_2017) and clustering tolerant to structural variability (vishwanathan_graph_2010).
Faults propagate only along explicit calls. Heavy co-location of MS instances (5.1) creates indirect propagation paths. Integrate call-graph edges with shared-infrastructure links in propagation models (boccaletti_multilayer_2014)
All MSs merit equally sophisticated scaling policies. Workload is highly skewed (5.2); most prediction benefit may lie in a small subset of MSs. Make resource-allocation and prediction policies total workload-aware.
Static dependency view is representative of the MSA. Temporal interactions can differ greatly from static topology (6.1). Combine temporal network representations (masuda_guide_2016) with static views.
MSA structure is stable enough that models trained once remain valid. MSAs evolve constantly (6.3), altering the normal baseline. Explore continual (parisi_continual_2019) and transfer learning (pan_transfer_2010) frameworks for online adaptation.
@@ 6.4. Dynamicity Implications
The static view of system dependencies does not necessarily correspond to runtime interactions: previous research has used the static view of dependencies to characterise MSAs (huye_lifting_2023) including identifying design anti-patterns (gamage_using_2021), assuming that the static view accurately represents MS behaviour at runtime. However, our results (Section 6.1) show that temporal interactions among MSs can diverge substantially from this static view, with marked differences in temporal centrality. Consequently, applying graph measures to the static dependency graph (gamage_using_2021) can yield characterisations that are misaligned with the system’s true runtime behaviour. Such analyses would benefit from complementing static representations with temporal network–based insights that capture runtime dynamics (masuda_guide_2016).
MS architectures evolve on a daily basis: of the commonly used MS testbeds (gan_open-source_2019; noauthor_fudanselabtrain-ticket_2025; noauthor_googlecloudplatformmicroservices-demo_2025; noauthor_ocp-power-demossock-shop-demo_2025) only one (noauthor_ciscodevnetbookinfo-cloudnative-sample_2025) incorporates versioning that changes the functionality of the deployment. This lack of evolving testbeds explains why most management approaches give limited attention to system evolution. A common general approach for management approaches is to model the “normal” behaviour of the system from historical data (yu_microrank_2021; gulenko_detecting_2018; wu_microdiag_2021; wu_microrca_2020; jin_anomaly_2020; xu_lightweight_2017; hacid_performance_2021; scheinert_learning_2021) or tracing data (wang_workflow-aware_2020; yang_transparently_2021; meng_detecting_2021; panahandeh_serviceanomaly_2024; nedelkoski_anomaly_2019; meng_midiag_2020; purfallah_mazraemolla_effective_2024), implicitly assuming a largely stationary architecture. Our findings (Section 6.3), consistent with Meta’s observations (huye_lifting_2023) and microservice design principles (newman_building_2021), show that MSAs change continuously: MSs and their dependencies appear, disappear, and are repurposed. Such changes inevitably alter call graph structure and, as MSs are multiplexed by new functionality, shift the expected patterns of their metrics. Although there are promising steps toward handling this non-stationarity, such as human-in-the-loop flagging of architectural changes (purfallah_mazraemolla_effective_2024) and transfer learning (gan_sleuth_2024), the practical implications of system evolution on management approaches, including the overhead of frequent retraining, remain under-appreciated.
7. Summary
@@
In Sections 4.5, 5.7, and 6.4 we examined how our findings challenge common assumptions in MSA management approaches. Table 4 summarises these findings together with their implications and recommendations for future work. Overall, our results reveal that existing approaches systematically underestimate the complexity of real-world deployments, particularly with respect to (i) scalability (Section 4.5), (ii) the fact that MSs may be multiplexed by many other MSs and front-end functionalities (Section 5.7), (iii) the variability of call graphs from the same front-end functionality (Section 5.7), (iv) the exposure of MSs to potential interference from co-habitation (Section 5.7), and (v) adaptation to an evolving system (Section 6.4).
@@
Many of the above limitations of current state-of-the-art approaches appear to stem, at least in part, from limited access to real-world MS deployments (huye_lifting_2023). The research testbeds that are commonly used throughout the literature (gan_open-source_2019; noauthor_fudanselabtrain-ticket_2025; noauthor_googlecloudplatformmicroservices-demo_2025; noauthor_ocp-power-demossock-shop-demo_2025) are not sophisticated enough to capture the complexities of actual deployments, especially large-scale deployments such as at Alibaba and Meta (huye_lifting_2023). Therefore, we join recent calls for improved testbed realism (seshagiri_sok_2022; huye_lifting_2023).
To bridge this gap, future testbeds should move beyond single-application deployments by supporting a broader range of front-end services (see Section 4.2). They should also capture the varying nature of real-world systems, such as enable architectural evolution through versioning (detti_ubench_2023) and integrating workload generators with real-world service usage data that can generate service-specific demand fluctuations (noauthor_locustio_nodate). Such variability is crucial for exposing changes in dependency structure, including the appearance and disappearance of call paths over time (Sections 4.3 and 5.6). Finally, to reproduce the real-world variability of call graphs, testbeds should incorporate mechanisms known to alter call paths in practice, such as graceful degradation under load or failure, cache-driven eviction policies, and conditional user inputs that trigger alternative execution branches.